OpenStudy (anonymous):

Faith mixes 14 liters of 12% acid solution with a 30% acid solution which results in an 18% acid solution. If x represents the number of liters of 30% acid solution used, which equation can be used to find the number of liters of 30% acid solution in the mixture? &Find the number of liters of 30% solution in the new mixture.

OpenStudy (anonymous):

#1 A) 1.68 + 0.3x = 0.18(x + 14) B) 2.52 + 0.3x = 0.12(x + 14) C) 4.2 + 0.18x = 0.3(x + 14) D) 1.68 + 0.18x = 0.3(x + 14) #2 A) 4.7 B) 21 C) 12 D) 7

OpenStudy (hoblos):

14(12) + 30x = (14+x)(18) =>x=7

OpenStudy (hoblos):

and it is A for the 1st part :D

OpenStudy (anonymous):

Thank youu!(:

OpenStudy (hoblos):

:)))))))))))

OpenStudy (anonymous):

i struggle a lot in Algebra so that helped a ton(:

OpenStudy (hoblos):

i'm here for you (: you just have to ask and i will answer ;)